OpenStudy (anonymous):

algebra 2 please help

OpenStudy (dumbcow):

got another one

OpenStudy (anonymous):

yes

OpenStudy (mathteacher1729):

@hinarauf80 -- what are you trying to do with the expression? (simplify, expand, round to the nth place, etc?)

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

oh wait sorry wrong question lolz simplify

OpenStudy (mathteacher1729):

@hinarauf80 -- what work have you done on this problem? Where are you getting stuck?

OpenStudy (dumbcow):

distribute

OpenStudy (dumbcow):

\[\sqrt{39}(\sqrt{6} +7) = \sqrt{39*6} +7\sqrt{39} = \sqrt{26*9}+7\sqrt{39} = 3\sqrt{26}+7\sqrt{39}\]